OpenStudy (anonymous):

a rectangle has a diagonal of 18.4cm and awidth of 7cm to the nearest centimeter what is the length of the rectangle doo ii*multiple the length and width ?

OpenStudy (anonymous):

use pythagoras \[a^2+b^2=c^2\] \[7^2+b^2=18.4^2 \] \[b^2=18.4^2-7^2\] get the right side with a calculator, then take the square root

OpenStudy (anonymous):

i get \[b^2=338.56-49=289.56\] \[b=\sqrt{289.56}=17\] rounded

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

welcome